Solve using substitution
y=3x+8
3y-x=7
I found the value of x, but I get stuck when trying to find y
y=3(-17/8)+8
Please help I know the answer x=13/8 but I want to know how to get the answer!!!
Thank you

Answer by ankor@dixie-net.com(22740)

Solve using substitution
y = 3x + 8
3y - x = 7
substitute (3x+8) for y
3(3x+8) - x = 7
9x + 24 - x = 7
9x - x = 7 - 24
8x = -17
x = -17/8, which you got
:
The using the 1st equation, replace x
y = 3(-17/8) + 8
y = -51/8 + 8
Change 8 to 64/8
y = -51/8 + 64/8
y = 13/8
:
:
Check solutions in the 2nd equation
3y - x = 7
3(13/8) - (-17/8) = 7
39/8 + 17/8 =
56/8 = 7